# Loss of integrability in a system with two-photon interactions

### v1

@@ IV.2.2 Large atomic level splitting: $\omega_{0}=2$
Having similar behaviors in the route to chaos observed in the average spectral ratio $\langle r\rangle$, there are noticeably differences between the two cases analyzed above. For the case with $\omega_{0}=\pi/5$, Figs. 4(a3)-4(c3) show a correspondence between the expectation values of $\hat{J}_{z}$ and the eigenvalues $m_{z}$, which is clearly absent for the case with $\omega_{0}=2$, shown in Figs. 5(a3)-5(c3).
@@ IV.3 Integrability and atomic level splitting
In Figs. 6(a1)–6(e1), we present the average spectral ratio $\langle r\rangle$ for increasing values of the atomic level splitting $\omega_{0}$, fixing the coupling strength at $\gamma=0.3$. The figures follow the same organization used previously, showing $\langle r\rangle$ alongside the Peres lattices for the operators $\hat{a}^{\dagger}\hat{a}$, $\hat{J}_{z}$, and $\hat{J}_{x}^{2}$. Each indicator is presented in Figs. 6(a1)–6(e1), Figs. 6(a2)–6(e2), Figs. 6(a3)–6(e3), and Figs. 6(a4)–6(e4), respectively.
Similar to the trends observed in the previous section with increasing coupling strength $\gamma$, we find that as the atomic level splitting $\omega_{0}$ increases, the system transitions from the integrable to the non-integrable regime. This is evident from the growing disorder in the Peres lattices as $\omega_{0}$ increases. The disorder is most prominent in the sequence of Peres lattices of the operator $\hat{J}_{x}^{2}$ [see Figs. 6(a4)–6(e4)]. However, even in cases with strong disorder, ordered patterns can still be identified in the low-energy regions, as shown in Figs. 6(c2)–6(c4), Figs. 6(d2)–6(d4), and Figs. 6(e2)–6(e4) in column view.
@@ IV.4 Onset of classical and quantum chaos
In this section, we investigate the classical dynamics of the two-photon Dicke model using the classical Hamiltonian given in Eq. (5). To explore the system’s behavior, we use Poincaré sections for various classical energy shells $\epsilon$, corresponding to regions where the quantum system exhibits either integrable or non-integrable behavior, as seen in the Peres lattices of the given operators. The second-order equation $h_{\text{D}}(q,p;Q,P)=\epsilon$ yields two roots for the variable $q=q_{\pm}(\epsilon,p,Q,P)$, and we select the positive root to define the initial conditions in the phase space, specifically setting $q_{0}=q_{+}(\epsilon,p_{0}=0,Q_{0},P_{0})$.
In Figs. 7(a1)-7(e1), we present Poincaré sections for an ensemble of initial conditions evolved in phase space using the two-photon Dicke Hamiltonian in Eq. (5), for increasing values of the classical energy shell $\epsilon$. These Poincaré sections are projected onto the atomic phase space defined by the classical variables $(Q,P)$. We fix the coupling strength at $\gamma=0.3$ and use a high atomic level splitting $\omega_{0}=2$, which corresponds to the parameter sets shown in Figs. 5(e1)-5(e4) and Figs. 6(e1)-6(e4) in column view. The chosen classical energy shells are within the region where we observe ordered patterns in the Peres lattices, transitioning to disorder as the average spectral ratio shifts from Poisson to GOE statistics. The last energy shells are displayed as vertical dashed lines in Figs. 6(e1)-6(e4) in column view. As seen in Figs. 7(a1)-7(e1), we detect an overall transition from integrability to chaotic motion with increasing energy. Specifically, the toroidal shapes in the Poincaré sections start to contract, and random patterns gradually emerge, indicating a shift toward complete ergodicity.
@@
Regarding the quantum-classical correspondence, we observe a reasonable agreement between the spectral statistics and the classical motion identified in the Poincaré sections. In Fig. 7(a1) and Fig. 7(a2), we see the correspondence between integrable classical motion and Poisson statistics. In contrast, Fig. 7(e1) and Fig. 7(e2) show the correspondence between fully chaotic classical motion and GOE statistics. Additionally, in Fig. 7(d1), we observe small stability islands surrounded by a chaotic sea, which do not significantly affect the spectral statistics described by the GOE distribution shown in Fig. 7(d2). For the other cases, we identify mixed classical dynamics, with corresponding spectral statistics that transition between Poisson and GOE distributions.
@@ V Conclusions
In this study, we carried out a comprehensive investigation into the loss of integrability in the two-photon Dicke model. We began by analyzing the integrable limit at zero atomic level splitting within the superradiant phase of the system, showing that the analytical solutions in this regime offer valuable insights for characterizing the breakdown of integrability in the low-energy spectrum. Building on this foundation, we shifted our focus to the normal phase, where we employed Peres lattices as a visual tool to track the onset of non-integrable behavior.
@@
To explore the classical dynamics, we used a mean-field approximation to derive the classical limit of the two-photon Dicke model in the normal phase. We confirmed the validity of this classical limit by establishing a quantum-classical correspondence between classical trajectories (via Poincaré sections) and quantum spectral statistics. In particular, we identified the onset of classical and quantum chaos in the high-energy regime of the system. Furthermore, we highlighted an intriguing regime of mixed dynamics Berry and Robnik (1984); Robnik (1998, 2023), which may offer a fertile ground for studying quantum localization phenomena Batistić and Robnik (2010, 2013a, 2013b). Of special interest is how this mixed classical behavior manifests in the quantum domain, where the coexistence of ordered and disordered structures in the Peres lattice coincides with a mixed spectral distribution [see Figs. 6(b1)-6(b4), column view].
Future directions include extending the analysis of classical dynamics into the superradiant phase and exploring whether classical and quantum chaos emerge in this regime. Additional avenues of research involve investigating quantum scarring, thermalization, and multifractality in the normal phase of the two-photon Dicke model, drawing parallels with known phenomena in the single-photon Dicke model Pilatowsky-Cameo et al. (2021); Villaseñor et al. (2023); Bastarrachea-Magnani et al. (2024).
@@ Acknowledgments
We acknowledge the support of the Computation Center - ICN, in particular to Enrique Palacios, Luciano Díaz, and Eduardo Murrieta. F.R., V.S.M-G, D.Y.C, and J.G.H acknowledge partial financial support from the DGAPA-UNAM project IN109523. D.V. was supported by the Slovenian Research and Innovation Agency (ARIS) under Grants No. J1-4387 and No. P1-0306.
@@ Appendix A Mean-field analysis with coherent states
A mean-field approximation is achieved by taking a trial state and computing the expectation value of an algebraic quantum Hamiltonian. The two-photon Dicke model is an algebraic Hamiltonian composed of two Hilbert subspaces, one bosonic $\mathcal{H}_{\text{B}}$ with infinite dimension, and one atomic $\mathcal{H}_{\text{A}}$ with finite dimension $12j+$ 1. Therefore, we use Glauber and Bloch coherent states of minimum uncertainty Arecchi et al. (1972); Zhang et al. (1990), given in Eq. (2) and Eq. (3), which are related to the bosonic and atomic subspace, respectively. By taking the expectation value of the two-photon Dicke Hamiltonian in Eq. (1) under tensor product of the last coherent states in Eq. (4), we get the expression
@@
Each coherent state parameter, $(\alpha,\beta)\in\mathbb{C}$, can be linked to a pair of canonical position and momentum variables in phase space Villaseñor et al. (2024)
@@
such that, $(q,p)$ identify the bosonic classical coordinates and $(Q,P)$ the atomic classical coordinates. By substituting the coherent state parameters $\alpha$ and $\beta$ in Eq. (17) and scaling the expectation value by the system size $j$, we get the classical two-photon Dicke Hamiltonian given in Eq. (5).
@@ Appendix B Bogoliubov transformation
The last Hamiltonian commutes with the operator $\hat{J}_{x}$, $[\hat{H}^{\prime}_{\text{D}},\hat{J}_{x}]=0$. Both operators share a set of eigenstates $|n;j,m_{x}\rangle=|n\rangle\otimes|j,m_{x}\rangle$, such that, $\hat{a}^{\dagger}\hat{a}|n;j,m_{x}\rangle=n|n;j,m_{x}\rangle$ and $\hat{J}_{x}|n;j,m_{x}\rangle=m_{x}|n;j,m_{x}\rangle$. Using the basis $|j,m_{x}\rangle$, we obtain $12j+$ 1 Hamiltonians for each eigenvalue $m_{x}$,
@@
which can be diagonalized employing a Bogoliubov transformation of the form Emary and Bishop (2002b)
@@
where $\sigma\in\mathbb{C}$ and the operators $\hat{c}$ and $\hat{c}^{\dagger}$ satisfy the Heisenberg-Weyl algebra with $[\hat{c},\hat{c}^{\dagger}]=\hat{\mathbb{I}}$.
@@ Appendix C Numerical diagonalization and convergence
The diagonalization of the two-photon Dicke Hamiltonian in Eq. (1) is performed using a basis formed by the tensor product of Fock states, $|n\rangle$, for the bosonic subspace and angular-momentum states, $|j,m_{z}\rangle$, for the atomic subspace
@@
where $c_{n,m_{z}}^{k}=\langle n;j,m_{z}|E_{k}\rangle$. We define the probability distribution of finding $n$ photons in the $k$-th eigenstate
@@
where $\delta$ is a small tolerance parameter. When this condition is satisfied, it guarantees that the wavefunction of the $k$-th. eigenstate is converged within the truncated Hilbert space and will remain unchanged upon further increasing $n_{\max}$. This criterion is checked for each eigenstate, ordered by ascending energy. The lowest energy eigenstate that violates this inequality determines the set of converged eigenstates (eigenenergies).

### v2

@@ IV.2.2 Nonresonant case: $\omega_{0}=(\pi/5)\omega$
Having similar behaviors in the route to chaos observed in the average spectral ratio $\langle r\rangle$, there are noticeably differences between the two cases analyzed above. For the nonresonant case with $\omega_{0}=(\pi/5)\omega$, Figs. 5(a3)-5(c3) show a correspondence between the expectation values of $\hat{J}_{z}$ and the eigenvalues $m_{z}$, which is clearly absent for the resonant case with $\omega_{0}=2\omega$, shown in Figs. 4(a3)-4(c3).
@@ IV.3 Integrability and atomic level splitting
In Figs. 6(a1)–6(e1), we present the average spectral ratio $\langle r\rangle$ for increasing values of the atomic level splitting $\omega_{0}$, fixing the coupling strength at $\gamma=0.3$. The figure follows the same organization used previously, showing $\langle r\rangle$ alongside the Peres lattices for the operators $\hat{a}^{\dagger}\hat{a}$, $\hat{J}_{z}$, and $\hat{J}_{x}^{2}$. Each indicator is presented in Figs. 6(a1)–6(e1),  6(a2)–6(e2),  6(a3)–6(e3), and 6(a4)–6(e4), respectively.
Similar to the trends observed in the previous section with increasing coupling strength $\gamma$, we find that as the atomic level splitting $\omega_{0}$ increases, the system transitions from the integrable to the nonintegrable regime. This is evident from the growing disorder in the Peres lattices as $\omega_{0}$ increases. The disorder is most prominent in the sequence of Peres lattices of the operator $\hat{J}_{x}^{2}$ [see Figs. 6(a4)–6(e4)]. However, even in cases with strong disorder, ordered patterns can still be identified in the low-energy regions, as shown in Figs. 6(c2)–6(c4),  6(d2)–6(d4), and 6(e2)–6(e4) in column view.
@@ IV.4 Onset of classical and quantum chaos
In this section, we investigate the classical dynamics of the two-photon Dicke model using the classical Hamiltonian given in Eq. (5). To explore the system’s behavior, we use Poincaré sections for various classical energy shells $\epsilon$, corresponding to regions where the quantum system exhibits either integrable or nonintegrable behavior, as seen in the Peres lattices of the given operators. The second-order equation $h_{\text{D}}(q,p;Q,P)=\epsilon$ yields two roots for the variable $q=q_{\pm}(\epsilon,p,Q,P)$, and we select the positive root to define the initial conditions in the phase space, specifically setting $q_{0}=q_{+}(\epsilon,p_{0}=0,Q_{0},P_{0})$.
In Figs. 7(a1)-7(e1), we present Poincaré sections for an ensemble of initial conditions evolved in phase space using the two-photon Dicke Hamiltonian in Eq. (5), for increasing values of the classical energy shell $\epsilon$. These Poincaré sections are projected onto the atomic phase space defined by the classical variables $(Q,P)$. We fix the coupling strength at $\gamma=0.3$ and use the resonant case of the atomic level splitting $\omega_{0}=2\omega$, which corresponds to the parameter sets shown in Figs. 4(e1)-4(e4) and Figs. 6(e1)-6(e4) in column view. The chosen classical energy shells are within the region where we observe ordered patterns in the Peres lattices, transitioning to disorder as the average spectral ratio shifts from Poisson to GOE statistics. The last energy shells are displayed as vertical dashed lines in Figs. 6(e1)-6(e4) in column view. As seen in Figs. 7(a1)-7(e1), we detect an overall transition from integrability to chaotic motion with increasing energy. Specifically, the toroidal shapes in the Poincaré sections start to contract, and random patterns gradually emerge, indicating a shift toward complete ergodicity.
@@
Regarding the quantum-classical correspondence, we observe a reasonable agreement between the spectral statistics and the classical motion identified in the Poincaré sections. In Figs. 7(a1) and 7(a2), we see the correspondence between integrable classical motion and Poisson statistics. In contrast, Figs. 7(e1) and 7(e2) show the correspondence between fully chaotic classical motion and GOE statistics. Additionally, in Fig. 7(d1), we observe small stability islands surrounded by a chaotic sea, which do not significantly affect the spectral statistics described by the GOE distribution shown in Fig. 7(d2). For the other cases, we identify mixed classical dynamics, with corresponding spectral statistics that transition between Poisson and GOE distributions.
@@ V Conclusions
In this study, we carried out a comprehensive investigation into the loss of integrability in the two-photon Dicke model. We began by analyzing the integrable limit at zero atomic level splitting within the superradiant phase of the system, showing that the analytical solutions in this regime offer valuable insights for characterizing the breakdown of integrability in the low-energy spectrum. Building on this foundation, we shifted our focus to the normal phase, where we employed Peres lattices as a visual tool to track the onset of nonintegrable behavior.
@@
To explore the classical dynamics, we used a mean-field approximation with coherent states to derive the classical limit of the two-photon Dicke model in the normal phase. We confirmed the validity of this classical limit by determining the ground-state energy and establishing a quantum-classical correspondence between classical trajectories (via Poincaré sections) and quantum spectral statistics. In particular, we identified the onset of classical and quantum chaos in the high-energy regime of the system. Furthermore, we highlighted an intriguing regime of mixed dynamics Berry and Robnik (1984); Robnik (1998, 2023), which may offer a fertile ground for studying quantum localization phenomena Batistić and Robnik (2010, 2013a, 2013b). Of special interest is how this mixed classical behavior manifests in the quantum domain, where the coexistence of ordered and disordered structures in the Peres lattice coincides with a mixed spectral distribution [see Figs. 6(b1)-6(b4), column view].
Future directions include extending the analysis of classical dynamics into the superradiant phase and exploring whether classical and quantum chaos emerge in this regime. Additional avenues of research involve investigating quantum scarring, thermalization, and multifractality in the normal phase of the two-photon Dicke model, drawing parallels with known phenomena in the single-photon Dicke model Pilatowsky-Cameo et al. (2021); Villaseñor et al. (2023); Bastarrachea-Magnani et al. (2024). Finally, it would be intriguing to explore the experimental verification of dynamical signatures of quantum chaos Das et al. (2025); Vallejo-Fabila et al. (2025) in this system.
@@ Acknowledgments
We acknowledge the support of the Computation Center - ICN, in particular to Enrique Palacios, Luciano Díaz, and Eduardo Murrieta. F.R., V.S.M-G, D.Y.C, and J.G.H acknowledge partial financial support from the DGAPA-UNAM Project No. IN109523. D.V. was supported by the Slovenian Research and Innovation Agency (ARIS) under Grants No. J1-4387 and No. P1-0306.
@@ Appendix A Mean-field analysis with coherent states
A mean-field approximation is achieved by taking a trial state and computing the expectation value of an algebraic quantum Hamiltonian. The two-photon Dicke model is an algebraic Hamiltonian composed of two Hilbert subspaces, one bosonic $\mathcal{H}_{\text{B}}$ with infinite dimension, and one atomic $\mathcal{H}_{\text{A}}$ with finite dimension $2j+1$. Therefore, we use Glauber and Bloch coherent states of minimum uncertainty Arecchi et al. (1972); Zhang et al. (1990), given in Eq. (2) and Eq. (3), which are related to the bosonic and atomic subspace, respectively. By taking the expectation value of the two-photon Dicke Hamiltonian in Eq. (1) under tensor product of the last coherent states in Eq. (4), we get the expression
@@
Each coherent-state parameter $(\alpha,\beta)\in\mathbb{C}$ can be linked to a pair of canonical position and momentum variables in phase space Villaseñor et al. (2024)
@@
such that, $(q,p)$ identify the bosonic classical coordinates and $(Q,P)$ the atomic classical coordinates. By substituting the coherent-state parameters $\alpha$ and $\beta$ in Eq. (18) and scaling the expectation value by the system size $j$, we get the classical two-photon Dicke Hamiltonian given in Eq. (5).
The minimization of the Hamilton equations of motion for the classical two-photon Dicke Hamiltonian yields the extreme values $\mathbf{x}_{\text{gs}}=(0,0;0,0)$, which correspond to the normal phase of the system. Evaluating Eq. (5) with these coordinates allows us to determine the ground-state energy for the normal phase
$\epsilon_{0}=\frac{E_{0}}{j}=h_{\text{D}}(\mathbf{x}_{\text{gs}})=-\omega_{0},$ (21)
which is the same expression obtained with the mean-field approximation using squeezed states Garbe et al. (2017).
@@ Appendix B Bogoliubov transformation
The last Hamiltonian commutes with the operator $\hat{J}_{x}$, $\hat{H}^{\prime}_{\text{D}},\hat{J}_{x}]=0$. Both operators share a set of eigenstates $|n;j,m_{x}\rangle=|n\rangle\otimes|j,m_{x}\rangle$, such that, $\hat{a}^{\dagger}\hat{a}|n;j,m_{x}\rangle=n|n;j,m_{x}\rangle$ and $\hat{J}_{x}|n;j,m_{x}\rangle=m_{x}|n;j,m_{x}\rangle$. Using the basis $|j,m_{x}\rangle$, we obtain $2j+1$ Hamiltonians for each eigenvalue $m_{x}$,
@@
which can be diagonalized by employing a Bogoliubov transformation of the form Emary and Bishop (2002b)
@@
where $\sigma\in\mathbb{C}$ and the operators $\hat{c}$ and $\hat{c}^{\dagger}$ satisfy the Heisenberg-Weyl algebra with $\hat{c},\hat{c}^{\dagger}]=\hat{\mathbb{I}}$.
@@
Figure 8: Fraction of converged eigenstates $N_{c}/N$ as a function of the coupling strength $\gamma$. The fraction of converged eigenstates was computed using the criterion of Eq. (36) with a tolerance parameter $\delta=10^{-3}$. In the first column, we set a system size $j=25$, two values of the atomic level splitting (a1) $\omega_{0}=0.0075$ and (a2) $\omega_{0}=2$, and three truncation values $n_{\max}=100,200$, and 500. In the second column, we set a truncation value $n_{\max}=500$, two values of the atomic level splitting (b1) $\omega_{0}=0.0075$ and (b2) $\omega_{0}=2$, and three system sizes $j=15,25$, and 40. In panel (a1), the vertical red dashed line represents the critical coupling strength $\gamma_{\text{c}}=\sqrt{\omega\omega_{0}j/2}$. In panel (b1), the vertical colored dashed lines represent the critical coupling strength for each system size $j=15,25$, and 40. In all panels, the vertical black dashed line represents the coupling strength of the spectral collapse $\gamma_{\text{sc}}=\omega/2$. The field frequency is $\omega=1$.
@@ Appendix C Numerical diagonalization and convergence
The diagonalization of the two-photon Dicke Hamiltonian in Eq. (1) is performed using a basis formed by the tensor product of Fock states $|n\rangle$ for the bosonic subspace and angular-momentum states $|j,m_{z}\rangle$ for the atomic subspace
@@
where $c_{n,m_{z}}^{k}=\langle n;j,m_{z}|E_{k}\rangle$. We define the probability distribution of finding $n$ photons in the $k$th eigenstate
@@
where $\delta$ is a small tolerance parameter. When this condition is satisfied, it guarantees that the wavefunction of the $k$th. eigenstate is converged within the truncated Hilbert space and will remain unchanged upon further increasing $n_{\max}$. This criterion is checked for each eigenstate, ordered by ascending energy. The lowest-energy eigenstate that violates this inequality determines the set of converged eigenstates (eigenenvalues).
In the following, we illustrate the effects of our convergence method. We first examine how the truncation value $n_{\max}$ impacts the fraction of converged eigenstates (eigenvalues). This fraction is defined as the ratio of the number of converged eigenstates $N_{c}$, identified using the criterion in Eq. (36), to the total number of eigenstates $N$.
In Fig. 8(a1), we consider an atomic level splitting of $\omega_{0}=0.0075$, where there is a coexistence of normal and superradiant phases before the spectral collapse. Figure 8(a2) examines a value of $\omega_{0}=2$, where only the normal phase exists prior to the spectral collapse. Both scenarios were computed for a fixed system size of $j=25$ and three truncation values $n_{\max}=100,200$, and 500. In Fig. 8(a1), we observe that the fraction $N_{c}/N$ increases slowly at the coupling strength corresponding to the spectral collapse $\gamma_{\text{sc}}=\omega/2$, particularly for the largest truncation value $n_{\max}=500$. For the other coupling strengths $\gamma<\omega/2$, the fraction remains relatively constant, regardless of the truncation value. In contrast, as shown in Fig. 8(a2), increasing the truncation value consistently enhances the fraction of converged states across all coupling strengths.
Finally, we examine the effect of system size $j$ on the fraction $N_{c}/N$. In Figs. 8(b1) and 8(b2), we analyze the previous values of atomic level splitting $\omega_{0}=0.0075$ and $\omega_{0}=2$, respectively. Both cases were computed with a fixed truncation value of $n_{\max}=500$ and for three different system sizes $j=15,25$, and 40. The results in Figs. 8(b1) and 8(b2) indicate that, for a constant truncation value, increasing the system size does not significantly alter the fraction $N_{c}/N$.